# Can we accurately read or write quantum data?

### v2

@@ Abstract
quantum measurement; no-go theorem; Wigner-Araki-Yanase theorem; quantum computing.
I Introduction
@@
Therefore, the possible limitations of quantum measurements are of particular interest. Wigner [35, 4], followed by Araki and Yanase [1], realized that additive conservation laws restrict what sharp observables can be accurately measured without being disturbed. Wigner showed that the ideal measurement of the spin along an axis is prevented by the conservation of the total angular momentum along an orthogonal axis. From this result, known as the Wigner-Araki-Yanase (WAY) theorem [17], it seems that we have to choose between inaccuracy and disturbing the system even when it is in an eigenstate of the observable.
@@
Even if the same measurement cannot be both accurate and non-disturbing, for some applications it seems enough to achieve accuracy and non-disturbance in separate measurements. We can write or prepare the states using non-disturbing measurements, and read the resulting states using accurate measurements.
While Wigner concluded in [4] that accurate but not repeatable (disturbing) measurements are possible even without the requirement that the apparatus is very large, Ohira and Pearle showed that this would violate the Yanase condition [24]. The Yanase condition [37] requires the probe observable to commute with the conserved quantity, and it is necessary if we want to observe the result of the measurement.
Despite Wigner’s well known interest in symmetries, he seemed to use the existence of a conservation law as a mere signature of unitarity. But there seems to be a deeper lesson about symmetry encoded in the WAY theorem. A deep relation between symmetries, measurements, and quantum reference frames was clarified by Loveridge, Miyadera, and Busch [18]. However, the constraints on measurements extend beyond symmetries, as we can see from Ozawa’s result that measurements of non-discrete observables are not repeatable [25], and from Tukiainen’s generalization of the WAY theorem beyond conservation laws [33]. For a review of results about limitations of quantum measurements, up to that date, see Busch [3]. The WAY theorem was later extended to the case when the conserved quantity has a continuous and unbounded spectrum [15]. The measurability and disturbance parts of the WAY theorem were also extended to unsharp observables (positive operator-valued measures), in the presence of bounded conserved quantities, and necessary conditions for either accurate (in an unsharp, probabilistic sense) or non-disturbing measurements can be achieved [22]. But the full scope of the WAY theorem and its generalizations is still being charted.
@@
Here I show that if the Hamiltonian is bounded from below, no quantum measurement of a sharp observable, and no state preparation, can be accurate. The main result is proved for pure states in Section §II, and generalized to mixed states in Section §III. Section §IV concludes the article with a brief discussion of the implications of these results.
II The main result
We consider quantum measurements of an observable represented by a self-adjoint operator $\mathsf{A}$. The observed system $S$ is assumed to be initially in a state represented by the unit vector $\psi(0)$ from a Hilbert space $\mathcal{H}_{S}$. The initial, “ready” state of the measuring device $M$ is denoted by a unit vector $\varphi_{\varnothing}$ from a Hilbert space $\mathcal{H}_{M}$.
Let the measured observable $\mathsf{A}$ be, in terms of projectors $\mathsf{P}_{\lambda}$ acting on $\mathcal{H}_{S}$ and adding up to the identity,
@@ II The main result
Let the pointer observable of the measuring device be, in terms of projectors $\Pi_{\xi}$ adding up to the identity,
@@
The combined system consisting of the measuring device and the observed system is considered isolated for the duration of the measurement. The evolution is governed by the unitary operator $\mathbf{U}_{t}=e^{-\frac{i}{\hbar}t\mathsf{H}}$. During a time interval $[0,T]$, the two systems interact with one another. We consider interactions of the form
$\mathbf{U}_{T}(\psi(0)\otimes\underbrace{\varphi_{\varnothing}}_{\mathclap{\in\Pi_{\varnothing}\mathcal{H}_{M}}})=\sum_{\lambda}c_{\lambda}\psi_{\lambda}(T)\otimes\underbrace{\varphi_{\lambda}(T)}_{\in\Pi_{\lambda}\mathcal{H}_{M}},$ (3)
where $\varphi_{\varnothing}$ and each $\varphi_{\lambda}(T)$ are eigenstates of the pointer observable $\mathsf{Z}$.
@@
An accurate measurement happens if $\mathbf{U}_{T}$ maps any separable state $\psi_{\lambda}(0)\otimes\varphi_{\varnothing}$, where $\psi_{\lambda}(0)$ is an eigenstate of $\mathsf{A}$, to another separable state $\psi_{\lambda}(T)\otimes\varphi_{\lambda}(T)$,
$\mathbf{U}_{T}\big{(}\underbrace{\psi_{\lambda}(0)}_{\in\mathsf{P}_{\lambda}\mathcal{H}_{S}}\otimes\varphi_{\varnothing}\big{)}=\psi_{\lambda}(T)\otimes\varphi_{\lambda}(T).$ (4)
In this case, from the pointer indicating the value $\lambda$ of the pointer observable $\mathsf{Z}$, we can infer that the observed system was, immediately before the measurement, in an eigenstate of $\mathsf{A}$ with eigenvalue $\lambda$.
@@
An accurate preparation happens if, from the pointer indicating the value $\lambda$ at the end of the measurement $t=T$, we can infer that $\psi_{\lambda}(T)$ is in an eigenstate of $\mathsf{A}$ with eigenvalue $\lambda$,
$\mathbf{U}_{T}\big{(}\psi_{\lambda}(0)\otimes\varphi_{\varnothing}\big{)}=\underbrace{\psi_{\lambda}(T)}_{\in\mathsf{P}_{\lambda}\mathcal{H}_{S}}\otimes\varphi_{\lambda}(T).$ (5)
In both cases, at time $t=T$ we retain from the superposition (3) a single outcome $\psi_{\lambda}\otimes\varphi_{\lambda}(T)$, by invoking our favorite recipe, for example the Projection Postulate or an alternative proposal based on decoherence like the many-worlds interpretation or the pilot-wave theory,
$\mathbf{U}_{T}\big{(}\psi_{\lambda}(0)\otimes\varphi_{\varnothing}\big{)}=\psi_{\lambda}(T)\otimes\varphi_{\lambda}(T)$ (6)
@@
Assumption 1 (Pointer stability).
There is a time interval $[T,T^{\prime}]$, $T^{\prime}>T$, so that, if the total system $S+M$ remains isolated for $t\in[T,T^{\prime}]$, the pointer states $\varphi_{\lambda}(t)$ remain eigenstates of $\mathsf{Z}$ for each eigenvalue $\lambda$, that is,
$|\varphi_{\lambda}(t)\rangle\in\Pi_{\lambda}\mathcal{H}_{M}$ (7)
for any $t\in[T,T^{\prime}]$ and any $\lambda$, where $\Pi_{\lambda}\mathcal{H}_{M}$ is the eigenspace of $\mathsf{Z}$ corresponding to the eigenvalue $\lambda$ of $\mathsf{Z}$, cf.  equation (2).
Remark 1 (Generality).
The description of a quantum measurement often involves an interaction Hamiltonian,
$\mathsf{H}=\mathsf{H}_{S}\otimes\mathsf{I}_{M}+\mathsf{I}_{S}\otimes\mathsf{H}_{M}+\mu\mathsf{H}_{S+M},$ (8)
where $\mathsf{H}_{S}$ and $\mathsf{H}_{M}$ are the free Hamiltonians of the observed system and the apparatus, and $\mathsf{I}_{S}$ and $\mathsf{I}_{M}$ are the identity operators on the two Hilbert spaces. The measurement involves turning on the coupling $\mu$ between the observed system and the measuring device, and after the measurement turning it off. This is sometimes assumed to be done by an external system, for example an agent, and makes the total Hamiltonian time-dependent. A description in terms of a time-independent Hamiltonian, as for an isolated system, requires the inclusion of this other system, in addition $S$ and $M$. We can add to our equations a Hilbert space $\mathcal{H}_{\mathcal{E}}$ of the environment. However, we can also consider that the Hilbert space $\mathcal{H}_{M}$ includes the additional system, and in fact the entire laboratory, considered isolated. The strength of the interaction also depends on the distance between the observed system and the apparatus, for example spin measurements involve magnetic fields, and the interaction depends on the distance and the orientation of the observed particle. Here we consider the entire laboratory, as well as the position degrees of freedom of its constituent particles, to be included in the system $M$, and also that the observed system $S$ includes its own position degrees of freedom, so that the entire measurement process is autonomous. The pointer observable is then understood as an observable of the entire system $M$. Therefore, there is no loss of generality in the formulation that we are using here. This formulation covers all kinds of sharp measurements, as long as the “ready” pointer state and the states corresponding to the outcomes are assumed to be pure states. It can be verified that all versions of measurements of sharp observables, as listed for example in [6], are particular cases of the description presented here. In Section §III we will extend the results to mixed states, again in the most general settings for sharp measurements.
One may imagine more general scenarios, in which the pointer indicates the state of the observed system at a different time than $t=0$ (measurements) and $t=T$ (preparations). All of these will be covered by the proof of the following result:
@@
We will need Lemma 1 from [11], p. 378:
@@
Let $\mathsf{H}$ be a self-adjoint operator on a Hilbert space $\mathcal{H}$, $\mathcal{H}^{\prime}$ a closed subspace of $\mathcal{H}$, and $\Psi\in\mathcal{H}$. If $\mathsf{H}$ is bounded from below and $e^{-\frac{i}{\hbar}t\mathsf{H}}\Psi\in\mathcal{H}^{\prime}$ for all values of $t$ in an open interval $(T,T^{\prime})$, then $e^{-\frac{i}{\hbar}t\mathsf{H}}\Psi\in\mathcal{H}^{\prime}$ for all $t\in\mathbb{R}$. ∎
@@
    $\Psi=\psi(0)\otimes\varphi_{\varnothing}$.
@@
Indeed, from Assumption 1, $\mathbf{U}_{t}\Psi\in\mathcal{H}^{\prime}$ for all $t\in(T,T^{\prime})$, so the hypothesis of Lemma 1 is satisfied. Then, its conclusion is that $\mathbf{U}_{t}\Psi\in\mathcal{H}^{\prime}$ for all $t\in\mathbb{R}$,
$\psi(t)\otimes\varphi_{\varnothing}\in\mathcal{H}_{S}\otimes\Pi_{\lambda}\mathcal{H}_{M}\ (\forall t\in\mathbb{R}).$ (9)
In particular, for $t=0$, since $\mathbf{U}_{0}\big{(}\psi(0)\otimes\varphi_{\varnothing}\big{)}=\psi(0)\otimes\varphi_{\varnothing}$,
$\psi(0)\otimes\varphi_{\varnothing}\in\mathcal{H}_{S}\otimes\Pi_{\lambda}\mathcal{H}_{M}.$ (10)
@@
$\varphi_{\varnothing}\in\Pi_{\lambda}\mathcal{H}_{M},$ (11)
in contradiction with the assumption that the “ready” pointer state $\varphi_{\varnothing}\in\Pi_{\varnothing}\mathcal{H}_{M}$, or, in general, that it cannot be an eigenstate of $\mathsf{Z}$ for the eigenvalue $\lambda$!
@@
As defined here, measurements ensure that if the observed system is in an eigenstate of $\mathsf{A}$ with eigenvalue $\lambda$ at $t=0$, the pointer indicates the value $\lambda$, so measurements amount to reading data about the observed system. In the case of preparations, the observed system is in such an eigenstate at the end of the measurement, $t=T$, so preparations amount to writing data in the observed system’s state.
But even if the interaction between the measuring device and the observed system is such that the observed system is in the state indicated by the pointer at a different time than $t=0$ and $t=T$, this this cannot avoid the inaccuracy, because Lemma 1 implies that, if the Hamiltonian is bounded from below, equation (11) from the proof of Theorem 1 has to be true.
Remark 3 (Relation with the Yanase condition).
We didn’t appeal to the Yanase condition, which was proposed to ensure the readability of the measurement’s result [1, 26]. This condition requires the apparatus’ part of the additive conserved observable from the WAY theorem to commute with the pointer observable. The reason is that Assumption 1 is more general. For a Hamiltonian as in equation (8), if the coupling becomes $\mu=0$ after the measurement, the Hamiltonian itself is an additive conserved quantity $\mathsf{H}=\mathsf{H}_{S}\otimes\mathsf{I}_{M}+\mathsf{I}_{S}\otimes\mathsf{H}_{M}$, and the Yanase condition becomes
$[\mathsf{Z},\mathsf{H}_{M}]=\mathsf{0}.$ (12)
Therefore, in the case from equation (8), the Yanase condition (12) follows, for a Hamiltonian bounded from below, from Assumption 1 and Lemma 1.
III Generalization to mixed states

### v3

@@ Abstract
Keywords: quantum measurement; no-go theorem; Wigner-Araki-Yanase theorem; quantum computing.
1 Introduction
@@
Therefore, the possible limitations of quantum measurements, and the possibility to accurately measure or prepare a quantum state, are of particular interest [^1] The term “accuracy” is the technical term used in the literature about quantum measurements, and it can be understood as the precision with which the obtained value corresponds to the actual eigenstate of the observed system. Another term is “repeatability”, which means that if the measurement is repeated, the same result obtains, or, in other words, the measurement doesn’t disturb the observed system [1].. Wigner [2, 3], followed by Araki and Yanase [4], realized that additive conservation laws restrict what sharp observables can be accurately measured without being disturbed. Wigner showed that the ideal measurement of the spin along an axis is prevented by the conservation of the total angular momentum along an orthogonal axis. From this result, known as the Wigner-Araki-Yanase (WAY) theorem [5], it seems that we have to choose between inaccuracy and disturbing the system even when it is in an eigenstate of the observable.
@@
Even if the same measurement cannot be both accurate and non-disturbing, for some applications it seems enough to achieve accuracy and non-disturbance in separate measurements. We can write (or prepare) the states using non-disturbing measurements, and read (or measure) the resulting states using accurate measurements.
While Wigner concluded in [3] that accurate but non-repeatable (non-disturbing) measurements are possible even without the requirement that the apparatus is very large, Ohira and Pearle showed that this would violate the Yanase condition [6]. The Yanase condition [7] requires the probe observable to commute with the conserved quantity, and it is necessary if we want to observe the result of the measurement.
Despite Wigner’s well known interest in symmetries, he seemed to use the existence of a conservation law as a mere signature of unitarity. But there seems to be a deeper lesson about symmetry encoded in the WAY theorem. A deep relation between symmetries, measurements, and quantum reference frames was clarified by Loveridge, Miyadera, and Busch [8]. However, the constraints on measurements extend beyond the direct applications of symmetries, as we can see from Ozawa’s result that measurements of non-discrete observables are not repeatable [9], and from Tukiainen’s generalization of the WAY theorem beyond conservation laws [10]. For an up to date review of results about limitations of quantum measurements see Busch [1]. The WAY theorem was later extended to the case when the conserved quantity has a continuous and unbounded spectrum [11]. The measurability and disturbance parts of the WAY theorem were also extended to unsharp observables (positive operator-valued measures), in the presence of bounded conserved quantities, and necessary conditions for either accurate (in an unsharp, probabilistic sense) or non-disturbing measurements can be achieved [12]. But the full scope of the WAY theorem and its generalizations is still being charted.
@@
Here I show that if the Hamiltonian is bounded from below, no quantum measurement of a sharp observable, and no state preparation, can be accurate. The main result is proved for pure states in Section §2, and generalized to mixed states in Section §4. Section §6 concludes the article with a brief discussion of the implications of these results.
2 The main result
We consider the quantum measurement of an observable represented by a self-adjoint operator $\mathsf{A}$. The observed system $S$ is assumed to be initially in a state represented by the unit vector $\psi(0)$ from a Hilbert space $\mathcal{H}_{S}$. The initial, “ready” state of the measuring device $M$ is denoted by a unit vector $\varphi_{\varnothing}$ from a Hilbert space $\mathcal{H}_{M}$.
Let the measured observable $\mathsf{A}$ be, in terms of projectors $\mathsf{P}_{\lambda}$ adding up to the identity on $\mathcal{H}_{S}$,
@@ 2 The main result
Let the pointer observable $\mathsf{Z}$ of the measuring device be, in terms of projectors $\Pi_{\xi}$ adding up to the identity on $\mathcal{H}_{M}$,
@@
We make the following assumption about the total Hamiltonian:
Assumption 1 (No negative energy).
The Hamiltonian of the combined system consisting of the measuring device and the observed system is bounded from below.
We also make the following assumption about the measurement:
Assumption 2 (Isolation).
The combined system consisting of the measuring device and the observed system is isolated for the duration of the measurement.
The evolution of the combined system $S+M$ is governed by the unitary operator $\mathbf{U}_{t}=e^{-\frac{i}{\hbar}t\mathsf{H}}$. We consider interactions of the form
$\mathbf{U}_{T}(\psi(0)\otimes\underbrace{\varphi_{\varnothing}(0)}_{\mathclap{\in\Pi_{\varnothing}\mathcal{H}_{M}}})=\sum_{\lambda}c_{\lambda}\psi_{\lambda}(T)\otimes\underbrace{\varphi_{\lambda}(T)}_{\in\Pi_{\lambda}\mathcal{H}_{M}},$ (3)
where $\varphi_{\varnothing}(0)$ and each $\varphi_{\lambda}(T)$ are eigenstates of the pointer observable $\mathsf{Z}$ [^2] If the operator $\mathsf{Z}$ is nondegenerate, it may seem redundant to include the dependency of time in the notation when we already labeled the eigenstates with the eigenvalue, so it may seem that we could just write, for example, $\varphi_{\lambda}$ instead of $\varphi_{\lambda}(T)$. But we do it anyway because we want to allow degeneracy for maximal generality, and because we want to indicate the time.. Here, $\psi_{\lambda}(T)$ is not necessarily an eigenvalue of $\mathsf{A}$, the label $\lambda$ being used for summation.
@@
An accurate measurement happens if, assuming that the pointer was in a “ready” state $\varphi_{\varnothing}(0)$ at $t=0$, $\mathbf{U}_{T}$ maps any separable state $\psi_{\lambda}(0)\otimes\varphi_{\varnothing}(0)$, where $\psi_{\lambda}(0)$ is an eigenstate of $\mathsf{A}$, to another separable state $\psi_{\lambda}(T)\otimes\varphi_{\lambda}(T)$,
$\mathbf{U}_{T}\big(\underbrace{\psi_{\lambda}(0)}_{\in\mathsf{P}_{\lambda}\mathcal{H}_{S}}\otimes\varphi_{\varnothing}(0)\big)\underset{\begin{subarray}{c}\phantom{O}\\ \Longleftrightarrow\end{subarray}}{=}\psi_{\lambda}(T)\otimes\underbrace{\varphi_{\lambda}(T)}_{\in\mathsf{\Pi}_{\lambda}\mathcal{H}_{M}},$ (4)
where $\varphi_{\lambda}(T)$ is an eigenstate of the pointer observable $\mathsf{Z}$, and, from the pointer indicating the value $\lambda$ of the pointer observable $\mathsf{Z}$, we can infer that the observed system was, at $t=0$, in an eigenstate of $\mathsf{A}$ with eigenvalue $\lambda$. This is not assumed to be true if the observed system was, initially, not in an eigenstate of the observable or if the pointer was not in a “ready” state at $t=0$.
@@
An accurate preparation happens if, assuming that the pointer was in a “ready” state $\varphi_{\varnothing}(0)$ at $t=0$, from the pointer indicating the value $\lambda$ at the end of the measurement $t=T$, we can infer that $\psi_{\lambda}(T)$ is an eigenstate of $\mathsf{A}$ with eigenvalue $\lambda$,
$\mathbf{U}_{T}\big(\psi(0)\otimes\varphi_{\varnothing}(0)\big)=\sum_{\lambda}c_{\lambda}\underbrace{\psi_{\lambda}(T)}_{\in\mathsf{P}_{\lambda}\mathcal{H}_{S}}\underset{\begin{subarray}{c}\phantom{O}\\ \Longleftarrow\end{subarray}}{\otimes}\underbrace{\varphi_{\lambda}(T)}_{\in\mathsf{\Pi}_{\lambda}\mathcal{H}_{M}}.$ (5)
At time $t=T$ we retain from the superposition (3) a single outcome $\psi_{\lambda}(T)\otimes\varphi_{\lambda}(T)$, by invoking our favorite recipe, for example the Projection Postulate or an alternative proposal based on decoherence like the many-worlds interpretation or the pilot-wave theory,
$\sum_{\lambda}c_{\lambda}\psi_{\lambda}(T)\otimes\varphi_{\lambda}(T)\mapsto\psi_{\lambda_{\textnormal{o}}}(T)\otimes\varphi_{\lambda_{\textnormal{o}}}(T),$ (6)
where $\lambda_{\textnormal{o}}$ is a particular outcome of the measurement or the preparation.
@@
Assumption 3 (Pointer persistence).
There is a time interval $^{\prime}], T>TT^{\prime}>T$, so that, if the total system $S+M$ remains isolated for $t\in[T,T^{\prime}]$, the pointer states $\varphi_{\lambda}(t)$ remain eigenstates of $\mathsf{Z}$ for each eigenvalue $\lambda$, that is,
$\varphi_{\lambda}(t)\in\Pi_{\lambda}\mathcal{H}_{M}$ (7)
for all $t\in[T,T^{\prime}]$ and any $\lambda$, where $\Pi_{\lambda}\mathcal{H}_{M}$ is the eigenspace of $\mathsf{Z}$ corresponding to the eigenvalue $\lambda$ of $\mathsf{Z}$, cf. equation (2).
Justification.
Note that $\mathsf{Z}$ can be degenerate (and it usually is highly degenerate, because the pointer is a macroscopic object), and $\varphi_{\lambda}(t)$ can change in time, but, for $t\in[T,T^{\prime}]$, $\varphi_{\lambda}(t)$ is assumed to remain confined to the eigenspace $\Pi_{\lambda}\mathcal{H}_{M}$ of $\mathsf{Z}$, otherwise it would be impossible to know the result of the measurement. Whatever kind of quantum measurement we may have in mind, the experimentalist needs to be able to read the result. The outcome is recorded temporarily somewhere, on a photographic plate, in the position of a pointer and so on. It would be unrealistic to assume that it is recorded forever, but at least for a finite time interval $T,T^{\prime}]$, the experimentalist should be able to read the result. And the result is recorded in the eigenvalue of the pointer’s state vector, which is an eigenvector of the pointer observable. This minimal requirement absolutely necessary for experiments to be possible is expressed in Assumption 3. ∎
The Definitions 1 and 2 allow treating the measurement and the preparation as particular cases of equation (3), where the pointer indicates the state of the observed system at time $t=0$ (measurements) and $t=T$ (preparations). Both of these cases will be covered by the proof of the following result:
@@
We will need the Lemma from [24], p. 378:
@@
Let $\mathsf{H}$ be a self-adjoint operator on a Hilbert space $\mathcal{H}$, $\mathcal{H}^{\prime}$ a closed subspace of $\mathcal{H}$, and $\Psi_{0}\in\mathcal{H}$. If $\mathsf{H}$ is bounded from below and $e^{-\frac{i}{\hbar}t\mathsf{H}}\Psi_{0}\in\mathcal{H}^{\prime}$ for all values of $t$ in an open interval $(T,T^{\prime})$, then $e^{-\frac{i}{\hbar}t\mathsf{H}}\Psi_{0}\in\mathcal{H}^{\prime}$ for all $t\in\mathbb{R}$. ∎
@@
    $\Psi_{0}=\mathbf{U}_{T}^{-1}\big(\psi_{\lambda}(T)\otimes\varphi_{\lambda}(T)\big)$, where $\varphi_{\lambda}(T)$ is an eigenstate of the pointer observable.
@@
The conditions (4), respectively (5), apply, in particular, to initial states that evolve into a definite value for the pointer without having to appeal to the projection from equation (6). In the case of a measurement, this happens if
$\Psi_{0}=\psi(0)\otimes\varphi_{\varnothing}(0),$ (8)
where $\psi(0)$ is an eigenvector $\psi_{\lambda}(0)$ of $\mathsf{A}$ for some eigenvalue $\lambda$. In the case of a preparation, this happens if $\Psi_{0}$ is an eigenvector of the operator resulting by evolving back to $t=0$ the operator $\mathsf{P}_{\lambda}\mathcal{H}_{S}\otimes\mathsf{\Pi}_{\lambda}\mathcal{H}_{M}$. In both cases, the pointer is assumed to start in a “ready” state. Since both measurements and preparations should work as well for these situations in which the evolution is strictly unitary, we can use them to derive a contradiction.
From Assumption 3, $\mathbf{U}_{t}\Psi_{0}\in\mathcal{H}^{\prime}$ for all $t\in(T,T^{\prime})$, so the hypothesis of Lemma 1 is satisfied. Then, its conclusion is that $\mathbf{U}_{t}\Psi_{0}\in\mathcal{H}^{\prime}$ for all $t\in\mathbb{R}$ and all eigenvalues $\lambda$. In particular, for $t=0$, since $\mathbf{U}_{0}\big(\Psi_{0}\big)=\Psi_{0}$, we obtain
$\Psi_{0}\in\mathcal{H}_{S}\otimes\Pi_{\lambda}\mathcal{H}_{M}.$ (9)
@@
$\varphi_{\varnothing}(0)\in\Pi_{\lambda}\mathcal{H}_{M},$ (10)
in contradiction with the assumption that the “ready” pointer state $\varphi_{\varnothing}(0)\in\Pi_{\varnothing}\mathcal{H}_{M}$, or, in general, that $\varphi_{\varnothing}(0)$ cannot be an eigenstate of $\mathsf{Z}$ for the eigenvalue $\lambda$!
@@
We will deal more with the general cases when the observed system and the measuring device are entangled or even mixed states in Section §4.
@@
The proof is based on extracting conditions for the “ready” pointer state from condition (4), respectively (5), applied to different values of $\lambda$. For the eigenstates, the evolution is unitary, without projection, at least between $t=0$ and $t=T^{\prime}$, because the observed system already is in an eigenstate of the observable. As seen in the proof, these conditions turn out to require, for each eigenvalue $\lambda$, that the “ready” pointer state has to be the pointer state corresponding to $\lambda$, which is a contradiction.
Since conditions (4) and (5) are assumed to be satisfied by the measuring device in standard Quantum Mechanics, they are assumed in all interpretations, even those that include the Projection Postulate, because this Postulate is invoked only if the measurement process leads to a superposition of different outcomes.
3 Methods, explanations, and physical interpretation
The method used in the proof relies on obtaining a contradiction from the assumption that the Hamiltonian is bounded from below, so it is a proof by reductio ad absurdum. Theorem 1 also assumes that the measurement takes place in isolation (Assumption 2), and that the value indicated by the pointer persists for at least a finite duration (Assumption 3). The model of measurement is the most general that assumes that the observed system is initially separated from the measuring device and that the observable is sharp. However, the condition of separability is dropped in the generalization from Theorem 2, Section §4.
The proof starts with the conditions (4) and respectively (5), sometimes called in the literature calibration conditions, [25, 1, 26]. These conditions have to be satisfied regardless of the interpretation of Quantum Mechanics, and don’t require projection or collapse, because the state is already an eigenstate. The proof shows that, if the observed system is in an eigenstate for the eigenvalue $\lambda$ of $\mathsf{A}$, the assumption that the Hamiltonian is bounded from below (Assumption 1), combined with the other assumptions, implies that the pointer’s “ready” state $\varphi_{\varnothing}(0)$ has to be an eigenvalue for the eigenvalue $\lambda$ for $\mathsf{Z}$. Consider two distinct observables $\lambda\neq\lambda^{\prime}$. Then, the proof of Theorem 1 shows that $\varphi_{\varnothing}(0)$ has to be an eigenvector for both $\lambda$ and $\lambda^{\prime}$. But this is possible only if $\lambda=\lambda^{\prime}$, in which case there is no measurement, or if $\varphi_{\varnothing}=0$, but this vector can’t represent a pointer state because it is not a unit vector. The pointer states $\varphi_{\varnothing}(0)$, $\varphi_{\lambda}$, and $\varphi_{\lambda^{\prime}}$ have to be mutually orthogonal, as in Figure 1.
[…]
Figure 1: The contradiction derived in the proof is that the pointer’s ready state $\varphi_{\varnothing}$ has to be an eigenstate for both $\lambda$ and $\lambda^{\prime}$, while in fact $\varphi_{\varnothing}$, $\varphi_{\lambda}$, and $\varphi_{\lambda^{\prime}}$ have to be mutually orthogonal.
Perhaps the least intuitive part of the proof is Lemma 1, so let’s give here its proof, following [24].
Proof of Lemma 1.
Let $\Phi$ be an arbitrary vector orthogonal to $\mathcal{H}^{\prime}$. Let $z=t+iy$, with $y\leq 0$. Then, because $\mathsf{H}$ is bounded from below, the function $g(z):=\langle\Phi|e^{-i\mathsf{H}z}\Psi\rangle$ is analytic for $y<0$ (see for example [27]). The function $g(z)$ is continuous for $y\leq 0$ and, because $\Psi(t)\in\mathcal{H}^{\prime}$ for $t\in(T,T^{\prime})$, $g(z)$ vanishes for $z=t\in(T,T^{\prime})$. For the next step of the proof we apply a theorem in complex analysis, called the Schwarz reflection principle (see for example [28, 27]), according to which if the function $g(z)$ defined on $\{z=t+iy\in\mathbb{C}|y\leq 0\}$ is analyting for $y<0$, continuous for $y\leq 0$, and real for $z=t\in(T,T^{\prime})$, it can be extended analytically on the entire complex plane $\mathbb{C}$ by making $g(t-iy):=g(t+iy)^{\ast}$ for all $y>0$. Since $g(z)$ is analytic and vanishes on $(T,T^{\prime})$, it follows that it vanishes everywhere, in particular for all $t\in\mathbb{R}$. This implies that $\Phi\perp\mathsf{U}_{t}$ for all $t\in\mathbb{R}$, so $e^{-\frac{i}{\hbar}t\mathsf{H}}\Psi_{0}\in\mathcal{H}^{\prime}$ for all $t\in\mathbb{R}$. ∎
4 Generalization to mixed states